how many 20% acid solution should be mixed with 6 oz of 30% acid solution to produce a mixture that is 28% acid?
could u make a chart because I like visuals. I want to be able to use the same concept with other similar mixture problems. Thanks.

Answer by stanbon(75887)

how many 20% acid solution should be mixed with 6 oz of 30% acid solution to produce a mixture that is 28% acid?
------
Sketch 3 jugs
Start jug has x oz of 20% solution
Second jug as 6 oz of 30% solution
Result jug has (x+6)oz of 28% solution
-------------------------
Equation:
acid + acid = acid
0.20x + 0.30*6 = 0.28(x+6)
--------
20x + 30*6 = 28x + 28*6
----
8x = 2*6
x = 12/8 = 3/2 oz (amt. of 20% to add)
-------------------
Cheers,
Stan H.
---------------